40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 9 11:41:09 2015, ericq, Update, VAC, N2 Pressure fell 
    Reply  Fri Oct 16 18:27:52 2015, Steve, Update, VAC, vacuum normal condition oct16Fpm2015.png
       Reply  Tue Oct 20 16:27:04 2015, ericq, Update, VAC, vacuum VME machines rebooted 
          Reply  Wed Oct 21 10:03:15 2015, Steve, Update, VAC, after running out of N2 1d_after_reboot.png
             Reply  Wed Oct 21 15:42:56 2015, Steve, Update, VAC, vacuum gauge check Gauges.png
                Reply  Tue Oct 27 10:08:12 2015, Steve, Update, VAC, CC vacuum gauges gauges_5y_vs_8d.png13.9yCC1&4.png
                   Reply  Mon Dec 7 11:17:51 2015, Steve, Update, VAC, noisy RGA scan at day 433 433dRGAscan.png
                      Reply  Tue Dec 8 15:40:48 2015, Steve, Update, VAC, glitchy RGA scan at day 434 lessGlichingToday.png
                         Reply  Wed Dec 9 11:25:55 2015, Steve, Update, VAC, normal RGA scan at day 435-436 glichesGone.pngd436.png
                            Reply  Fri Jan 29 09:59:23 2016, Steve, Update, VAC, vacuum control upgrade 1100days.png
             Reply  Thu Oct 22 08:52:04 2015, Steve, Update, VAC, clean RGA scan after sweaty Maglev clean_scan.png
          Reply  Fri Sep 30 10:37:57 2016, Steve, Update, VAC, vacuum VME machines rebooted rebooted.png
Message ID: 12525     Entry time: Fri Sep 30 10:37:57 2016     In reply to: 11703
Author: Steve 
Type: Update 
Category: VAC 
Subject: vacuum VME machines rebooted 

[ Gautam and Steve ]

c1vac1 and c1vac2 were rebooted and the gauges are communicating now. V1, VA6, V5 and V4 were closed and disconnected to avoid unexpected valve switching. All went smoothly.

The new ITcc gauge is at 1e-5 Torr as CC1   This is the gauge that should be logged in slow channel.

TP2 fore line dry pump was replaced this morning after 382 day of operation.

TP3 dry pump is very noisy, but it's pressure still 47 mTorr

Quote:

[ericq, Gautam, Steve]

Following roughly the same procedure as ELOG 11354, c1vac1 and c1vac2 were rebooted. The symptoms were identical to the situation in that ELOG; c1vac1 could be pinged and telneted to, but c1vac2 was totally unresponsive. 

The only change in the linked procedure was that we did not shut down the maglev. Since I unwittingly had it running for days without V4 open while Steve was away, we now know that it can handle shorter periods of time than that...

Upon reboot, many channels were readable again, unfortunately the channels for TP2 and TP3 are still blank. We were able to return to "Vacuum normal state," but because of unknowned communication problems with VM1's interlock, we can't open VM1 for the RGA. Instead we opened VM2 to expose the RGA to the main IFO volumn, but this isn't part of the "Normal" state definite, so things currently read "Undefined state".

 

Attachment 1: rebooted.png  33 kB  Uploaded Fri Sep 30 11:40:27 2016  | Hide | Hide all
rebooted.png
ELOG V3.1.3-